On January 1, 2020, Buffalo Corporation purchased 350 of the $1,000 face value, 15%, 10-year bonds of Walters Inc. The bonds mature on January 1, 2030, and pay interest annually beginning January 1, 2021. Buffalo purchased the bonds to yield 11%. How much did Buffalo pay for the bonds?
Answer #1

Solution: Calculation of present value of the bond

Year Particulars Amount Present Value Factor Present Value
1-10 Interest $1000*350*15% = $52,500 5.8892 (11%,10 years) $52500 * 5.8892 = $309,183
10 Redemption Value $350,000 0.3522 $350,000 * 0.3522 = $123,270
Total present value $432,453

Buffalo corporation should pay $432,453 for purchasing of bonds.

Please rate positive and comment in case of any doubt. I would be happy to help you further.